# dc-GAN: Dual-Conditioned GAN for Face Demorphing From a Single Morph

### v1

@@ Abstract
A facial morph is an image created by combining two face images pertaining to two distinct identities. Face demorphing inverts the process and tries to recover the original images constituting a facial morph. While morph attack detection (MAD) techniques can be used to flag morph images, they do not divulge any visual information about the faces used to create them. Demorphing helps address this problem. Existing demorphing techniques are either very restrictive (assume identities during testing) or produce feeble outputs (both outputs look very similar). In this paper, we overcome these issues by proposing dc-GAN, a novel GAN-based demorphing method conditioned on the morph images. Our method overcomes morph-replication and produces high quality reconstructions of the bonafide images used to create the morphs. Moreover, our method is highly generalizable across demorphing paradigms (differential/reference-free). We conduct experiments on AMSL, FRLL-Morphs and MorDiff datasets to showcase the efficacy of our method. Code and train-test split is available at hidden for blind review.
Figure 1: An example of recovering the original bonafide images using our reference-free method is shown. Our method produces visually compelling reconstructions of constituent face images compared to previous approaches in Scenario 1. The generated outputs are i) distinct from each other and ii) noticeably different from the original morph.
@@ 1 Introduction
Figure 2: Dual-Conditioned GAN for Reference-Free Demorphing: An image encoder $\mathcal{E}$ encodes the morph image, which is then used as a condition in the generator. The generator, based on a UNet architecture $\mathcal{G}$, takes in the MORPH image and the encoded representation $\mathcal{E}(MORPH)$, producing two outputs, OUT1 and OUT2. The discriminator is trained to distinguish between the real set (MORPH, BF1, BF2) and the synthetic set (MORPH, OUT1, OUT2), differentiating real from synthetic pairs.
@@
Morphing, in the biometric literature, refers to the blending of two or more face images to generate a composite image that matches the identities if the images used to create it. This matching is with respect to an automated Face Recognition System (FRS). Such a composite image, known as a facial morph, can be incorporated into a biometric-based identity document thereby allowing multiple individuals to use the same document. This can compromise national security [25, 23]. In October 2018, German activists obtained a legitimate German passport by using a morphed picture of a European Union official and a member of their group[41]. Historically, morphs were created by extracting facial landmarks from face images and then combining the two face images [1, 14, 31]. However, more recently, deep learning models have been remarkably successful in generating high-quality morphs, especially, those based on generative paradigms like diffusion models [7] and generative adversarial attacks [12, 4, 8].
@@
Morph Attack Detection (MAD) is essential for retaining the integrity, security and reliability of FRS. Current MAD techniques can be broadly divided into two categories: i) reference-based differential-image technique [13, 34], and ii) reference-free single-image techniques [38, 5, 2]. While reference-based techniques compare the input image to a reference image, which is usually a trusted live image of the person taken under a trusted acquisition scenario, reference-free methods use facial features extracted from the input to determine whether or not the input is morphed. While MAD techniques can reliably flag facial morphs, they do not reveal the images involved in the creation of the morph. From a forensic standpoint, it is valuable to recover images of the constituent identities [^1] The constituent images are referred to as “bonafides” since in some literature, they are real images of real people. from a morph flagged by an MAD software. This process of extracting constituent face images from a morph image is called face demorphing. Limited work exists in this field, especially in reference-free face demorphing. In this paper, we focus on both differential as well as reference-free face demorphing.
@@
To set up the problem, we first identify multiple scenarios that are plausible in the real world. In the context of this work, the choice of train and test data will define these scenarios. To formally define the various scenarios arising in real life, let $\mathcal{X}_{train}$, $\mathcal{Y}_{train}$ denote the training data corresponding to the morphs and bonafides, respectively. Assume that all morph images in $\mathcal{X}_{train}$ images are necessarily generated only using the bonafide identities in $\mathcal{Y}_{train}$.
@@
where, $\mathcal{M}$ is the morphing operator. Test sets $\mathcal{X}_{test}$, $\mathcal{Y}_{test}$ are defined similarly. We now define three scenarios: i) Train and test morphs are obtained from the same pool of bonafide identities. ii) Some train and test morphs share the same bonafide identities but not all. iii) Train and test morphs are generated from a disjoint pool of bonafide identities. These conditions can be formally defined as:
@@
Most of the previous work assumed scenario 1 implicitly [38, 39], but this is a strong assumption hampering the generalisability of the demorphing technique. Limited research has been conducted on demorphing in Scenario 3 [5], and the results often replicate the morph in both the outputs (see Figure 1).
In this work, we propose a demorphing technique free of any assumption regarding the identities of images (scenario 3) as well as the availability of a reference image (differential/reference-free).
@@
  • We propose dc-GAN, a novel GAN-based technique designed to extract identities from a single morph. Our approach is well-suited for both reference-free and differential demorphing tasks.
  • We conduct a thorough analysis of the implicit bias present in morphs and extensively benchmark publicly available face recognition systems (FRS) on demorphing tasks in the presence of visible artefacts.
  • To the best of our knowledge, our method is the first to overcome morph-replication on unseen identities (scenario 3).
@@ 2 Background
Face Morphing refers to the process of combining two or more face images corresponding to two distinct identities to produce another face image (called a morph) which has a high biometric similarity (or less biometric distance), as assessed by an automated face matcher, with both identities. Typically, a complex morphing operator, $\mathcal{M}$, acts upon the input bonafides $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ to produce the morph $\mathcal{X}$:
@@
Face demorphing is the inverse of the morphing process. Given a morph $\mathcal{X}$, the goal is to recover the face images used to produce $\mathcal{X}$. Initial work on demorphing were primarily reference-based, i.e., they required an image of one of the identities to recover the other [13, 6]. The authors also assumed prior knowledge about the morphing factor used during the morphing process. FD-GAN [28] is a reference-based method which uses a dual architecture and attempts to recover an image if the first identity from the morphed input using an image of the second identity. To validate the effectiveness of its generative model, it then tries to recover the second identity using the output of the first identity. More recently, reference-free demporphing techniques have also been proposed with some degree of success. In [5], the authors decompose the morph image into two output images using a GAN that is composed of a generator, a decomposition critic, and two Markovian discriminators. Shukla [38] proposed a diffusion-based method for scenario 1 that iteratively adds noise to the morph image and recovers the constituent face images during the backward process. In [39], the authors decompose the morph into multiple privacy-preserving components using an encoder, and then recover the constituent face images by weighting and combining these components with a decoder.
@@ 3 Proposed Method
In [5], the authors attempt to recover identities from a single image using a GAN [15] that includes a generator for demorphing and three Markovian discriminators. However, the method has a problem, where the outputs not only closely resemble the morph image but also appear similar to each other (See Figure 1), we term this phenomenon as morph-replication. This issue arises because the morphing operator $\mathcal{M}$ is a highly nonlinear and complex function, making demorphing an ill-posed problem. Although their method employs three discriminators to optimize output similarity and visual realism, the generator only receives the morph image, which provides insufficient guidance for the demorphing process. The morph image alone may not be adequate to recover both ground-truth face images due to the biases introduced during the morphing process.
@@
For differential demorphing, we modify our pipeline slightly. The image encoder now encodes both the morph image and the reference image, concatenating them to provide guidance for the generator. The generator receives both the morph and reference images as input (with 6 channels) and outputs the reconstructed image (with 3 channels corresponding to RGB). The discriminator architecture and function remain unchanged.
@@ 3.2 Implementation details
Face Recognition Models and Parameters: We thoroughly evaluate our demorphing method using five publicly available Face Recognition Models: AdaFace [20], ArcFace [11], VGG-Face [26], FaceNet [35], and OpenFace [3]. We evaluate the performance of each FRS in the presence of high-frequency artefacts introduced during the generation process. While we benchmark five different FRSs, we use ArcFace as the primary FRS, consistent with its use in [5]. Each FRS is employed with its default settings as provided by the authors. We use cosine similarity to compute the biometric match score. Training is conducted using Adam optimization [21] with multi-GPU support via accelerate [16]. The training parameters are as follows: number of epochs: 300, learning rate: $10^{-4}$, dropout rate=0.1, $\beta_{1}$=0.5 and $\beta_{2}$=0.999.
@@ 5.1 Evaluation Criteria
Table 2: Reference Free Demorphing: We compute the True Match rate (TMR) at various False Match rate (FMR) thresholds (0.1/1/5/10) between the generated outputs (OUT1, OUT2) and the ground truth images (BF1, BF2) across the five FR models used. ArcFace performs most consistently across the FRS used in the presence of high-frequency artefacts (See Figure 5 and 6)
@@
Table 3: Differential Demorphing: We compute the True Match rate (TMR) at various False Match rate (FMR) thresholds between the generated output (OUT) and the ground truth image (GT) across the five FR models used.
@@
Figure 5: Reference-Free Demorphing Visualization: We illustrate the results of our reference-free demorphing approach. The model takes the morph image (MORPH) as input and produces two outputs (OUT1 and OUT2). The ground truth identities used to generate the morph are denoted as BF1 and BF2.
@@ 5.2.1 Reference Free Demorphing
We visualize the outputs of our reference-free demorphing method across six morphing techniques in Figure 5. These techniques include both landmark-based and deep-learning-based morphs. The first three columns visualize the morph and the two constituent face images used to create it, while the subsequent two columns show the outputs of our method, presented in no specific order. We observe that the faces produced by our method are visually distinct from each other and also appear notably different from the original morph. Moreover, they are visually closer to their ground-truth counterparts (refer to Appendix for more images). To quantify this, we calculate the normalized biometric distance between the generated faces and between the generated faces and the ground truth. We use five well-known face biometric models and plot the distance scores across three different pairings of images: OUT1-OUT2, BF1-OUT1, and BF2-OUT2, which represent the dissimilarity between the generated outputs and between each generated output and its corresponding ground truth. We plot the distance distribution for all datasets across all FR models in Figure 7. We observe an overall OUT1-OUT2 biometric distance of 0.4380, indicating a quantifiable difference in the output faces. Note that we use cosine distance rather than similarity to quantify results. This choice helps us intuitively explain morph-replication because a greater distance indicates that the generated faces are more distinguishable from each other. For BF1-OUT1, we observe a distance of 0.3992, and for BF2-OUT2, a distance of 0.4025. Note that the distance scores between outputs and their respective ground-truth face image fall well below the threshold used in several biometric systems [37, 36]. Our method successfully learns to distinguish features from the base image, such as hair (Rows 1 and 3), skin color, and subtle details like eyebrows and facial hair. Finally, we compute the True Match Rate (TMR) of the output subjects at various thresholds of False Match Rate (FMR) and present the results in Table 2. Our method achieves an average TMR of 93.86% on the AMSL dataset. The corresponding TMR values on the OpenCV, StyleGAN, WebMorph, FaceMorph, and MorDiff datasets are 93.99%, 58.92%, 89.87%, 94.39%, and 93.75%, respectively. We also evaluate the quality of images generated by our method using various image-quality assessment (IQA) metrics, as shown in Table 5. Our method achieves an average FID [17] of 0.4284 across all datasets, with corresponding averages of 0.61 for SSIM [42], 0.24 for LPIPS [44], and 18.02 for PSNR, demonstrating the visual realism of the outputs from our method. We also observe that ArcFace performs consistently across all datasets in the presence of visual artefacts compared to other FRS which is consistent with [5].
@@
Effect of $\mathcal{L}_{cr}$ and $\mathcal{E}(\cdot)$: We conduct an ablation study on our method to assess i) the impact of cross-road loss and ii) the influence of dual conditioning on generation. Excluding pixel-wise guidance (the second part of the loss in Equation 6) results in an average performance decline of 20.87% across the six datasets tested on the ArcFace FRS. This decrease is mainly due to the model’s reduced ability to accurately align the two unordered generated outputs with the ordered ground-truth images. Moreover, when the generator is trained without $\mathcal{E}$, the performance drops by 8.57% due to lack of context during the generation (see Section 3).
AMSL L 823 302 56 33
StyleGAN D 480 161 60 38
WebMorph L 480 160 60 37
FaceMorpher L 249 100 55 31
MorDiff D 195 62 43 26
OpenCV L 474 160 60 37

### v2

@@ Abstract
A facial morph is an image strategically created by combining two face images pertaining to two distinct identities. The goal is to create a face image that can be matched to two different identities by a face matcher. Face demorphing inverts the process and tries to recover the original images constituting a facial morph. Existing demorphing techniques are either very restrictive (assume identities during testing) or produce feeble outputs (both outputs look very similar - the morph-replication problem). In this paper, we overcome these issues by proposing dc-GAN (dual-conditioned GAN), a novel demorphing method conditioned on the morph image as well as the embedding extracted from the image. Our method overcomes morph-replication and produces high quality reconstructions of the bonafide images used to create the morphs. Moreover, our method is highly generalizable across demorphing paradigms (differential/reference-free). We conduct experiments on AMSL, FRLL-Morphs and MorDiff datasets to showcase the efficacy of our method. Code and train-test split are available at hidden for blind review.
@@ 1 Introduction
Morphing, in the biometric literature, refers to the blending of two or more face images to generate a composite image that matches the identities of the images used to create it. This matching is with respect to an automated Face Recognition System (FRS). Such a composite image, known as a facial morph, can be incorporated into a biometric-based identity document thereby allowing multiple individuals to use the same document. This can compromise national security [25, 23]. Historically, morphs were created by extracting facial landmarks from face images and then combining the two face images [1, 14, 31]. However, more recently, deep learning models have been remarkably successful in generating high-quality morphs, especially, those based on generative paradigms like diffusion models [7] and generative adversarial attacks [12, 4, 8].
@@
Morph Attack Detection (MAD) is essential for retaining the integrity, security and reliability of a biometric system. Current MAD techniques can be broadly divided into two categories: i) reference-based differential-image technique [13, 34], and ii) reference-free single-image techniques [38, 5, 2]. Reference-based methods have two images: the input face image (say, on the passport) and a reference face image (say, the live image of a person acquired at the port of entry). Reference-free methods have only one image: the input face image. While MAD techniques can reliably flag facial morphs, they do not reveal the images involved in the creation of the morph. From a forensic standpoint, it is valuable to recover images of the constituent identities [^1] The constituent images are referred to as “bonafides” since in some literature, they are real images of real people. from a morph flagged by an MAD software. This process of extracting constituent face images from a morph image is called face demorphing. Limited work exists in this field, especially in the reference-free category. In this paper, we focus on both reference-based as well as reference-free face demorphing.
Figure 1: Dual-Conditioned GAN for Reference-Free Demorphing: An image encoder, $\mathcal{E}$, encodes the morph image, which is then used as a condition in the generator. The generator, based on a UNet architecture, $\mathcal{G}$, takes in the MORPH image and the encoded representation, $\mathcal{E}(MORPH)$, producing two outputs, OUT1 and OUT2. The discriminator is trained to distinguish between the real set (MORPH, BF1, BF2) and the synthetic set (MORPH, OUT1, OUT2), differentiating real from synthetic pairs.
@@
To set up the problem, we first identify multiple scenarios that are plausible in the real world. In the context of this work, the choice of train and test data will define these scenarios. To formally define the various scenarios arising in real life, let $\mathcal{X}_{train}$, $\mathcal{Y}_{train}$ denote the training data corresponding to the morphs and bonafides, respectively. Assume that all morph images in $\mathcal{X}_{train}$ images are necessarily generated only using the identities in $\mathcal{Y}_{train}$.
@@
where, $\mathcal{M}$ is the morphing operator. Test sets $\mathcal{X}_{test}$, $\mathcal{Y}_{test}$ are defined similarly. We now define three scenarios: i) Train and test morphs are obtained from the same pool of identities. ii) Some train and test morphs share the same identities but not all. iii) Train and test morphs are generated from a disjoint pool of identities. These conditions can be formally defined as:
@@
Most of the previous work assumed scenario 1 implicitly [38, 39], but this is a strong assumption hampering the generalisability of the demorphing technique. Limited research has been conducted on demorphing in scenario 3 , and most techniques seem to merely replicate the morph in both the outputs[5].
In this work, we propose a demorphing technique free of any assumption regarding the identities of images (scenario 3) and can be used for both reference-based and reference-free demorphing. .
@@
  • We propose dc-GAN, a novel GAN-based technique designed to extract constituent images from a single morph. Our approach is well-suited for both reference-free and differential demorphing tasks.
  • We conduct a thorough analysis of the implicit bias present in morphs and extensively benchmark publicly available face recognition systems (FRS) on demorphing tasks in the presence of high frequency visible artefacts which arise during demorphing process.
  • To the best of our knowledge, our method is the first to overcome the issue of morph-replication on unseen identities (scenario 3).
@@ 2 Background
Face morphing refers to the process of combining two or more face images corresponding to two distinct identities to produce another face image (called a morph) which has a high biometric similarity (or lower distance), as assessed by an automated face matcher. Typically, a morphing operator, $\mathcal{M}$, acts upon two input bonafides, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, to produce the morph $\mathcal{X}$:
@@
Face demorphing is the inverse of the morphing process. Given a morph $\mathcal{X}$, the goal is to recover the face images used to produce $\mathcal{X}$. Initial work on demorphing were primarily reference-based, i.e., they required an image of one of the identities to recover the other [13, 6]. The authors also assumed prior knowledge about the morphing factor used during the morphing process. FD-GAN [28] is a reference-based method which uses a symmetric GAN architecture and attempts to recover an image of the first identity from the morphed input using an image of the second identity. To validate the effectiveness of its generative model, it then tries to recover the second identity using the output of the first identity. More recently, reference-free demporphing techniques have also been proposed with some degree of success. In [5], the authors decompose the morph image into two output images using a GAN that is composed of a generator, a decomposition critic, and two Markovian discriminators. Shukla [38] proposed a diffusion-based method for scenario 1 that iteratively adds noise to the morph image and recovers the constituent face images during the backward process. In [39], the authors decompose the morph into multiple privacy-preserving components using an encoder, and then recover the constituent face images by weighting and combining these components with a decoder.
@@ 3 Proposed Method
Face demorphing is the inverse of the morphing problem. The intricacy lies not only in the absence of the morphing technique used (landmark-based or deep learning based) but also in the lack of constraints on the output space. Given a morph $\mathcal{X}$, the goal of the demorphing operator, denoted as $\mathcal{DM}$ is to recover the constituent face images,
$O_{1},O_{2}=\mathcal{DM}(\mathcal{X})$ (4)
satisfying the following conditions:
$||\mathcal{B}(O_{1})-\mathcal{B}(O_{2})||^{2}<\theta$ (5)
and,
$\max_{j\in\{1,2\}}\max_{\begin{subarray}{c}k\in\{1,2\}\\ k\neq j\end{subarray}}(||\mathcal{B}(O_{j})-\mathcal{B}(I_{k})||^{2},||\mathcal{B}(O_{j})-\mathcal{B}(I_{j})||^{2})>\epsilon$ (6)
where $\theta$ and $\epsilon$ are similarity thresholds. Equation 5 enforces the reconstructed outputs to look dissimilar among themselves, an issue referred to as morph replication, while Equation 6 enforces that each output aligns with its corresponding ground truth image.
In [5], the authors attempt to recover identities from a single image using a GAN [15] that includes a generator for demorphing and three Markovian discriminators. However, the method has a limitation, where the outputs not only closely resemble the morph image but also appear similar to each other. We term this phenomenon as morph-replication. This issue arises because the morphing operator $\mathcal{M}$ is a highly nonlinear function, making demorphing an ill-posed problem. Although their method employs three discriminators to optimize output similarity and visual realism, the generator only receives the morph image, which provides insufficient guidance for the demorphing process. The morph image alone may not be adequate to recover both constituent face images.
@@
For differential, i.e., reference-based demorphing, we modify our pipeline slightly. The image encoder now encodes both the morph image and the reference image, concatenating them to provide guidance for the generator. The generator receives both the morph and reference images as input (with 6 channels) and outputs the reconstructed image (with 3 channels corresponding to RGB). The discriminator architecture and function remain unchanged.
@@ 3.1 Loss Functions
AMSL L 823 302 56 33
StyleGAN D 480 161 60 38
WebMorph L 480 160 60 37
FaceMorpher L 249 100 55 31
MorDiff D 195 62 43 26
OpenCV L 474 160 60 37
@@ 3.2 Implementation details
Face Recognition Models and Parameters: We thoroughly evaluate our demorphing method using five publicly available Face Recognition Models: AdaFace [20], ArcFace [11], VGG-Face [26], FaceNet [35], and OpenFace [3]. We evaluate the performance of each FRS in the presence of high-frequency artefacts introduced during the generation process. While we benchmark five different FRS, we use ArcFace as the primary FRS, consistent with its use in [5]. Each FRS is employed with its default settings as provided by the authors. We use cosine similarity to compute the biometric match score. Training is conducted using Adam optimization [21] with multi-GPU support via accelerate [16]. The training parameters are as follows: number of epochs: 300, learning rate: $10^{-4}$, dropout rate=0.1, $\beta_{1}$=0.5 and $\beta_{2}$=0.999.
@@ 5.1 Evaluation Criteria
Table 3: Reference Free Demorphing: We compute the True Match Rate (TMR in %) at various False Match Rate (FMR in %) thresholds (0.1/1/5/10) between the generated outputs (OUT1, OUT2) and the ground truth images (BF1, BF2) across the five FR models used. ArcFace performs most consistently across the FRS used in the presence of high-frequency artefacts (See Figure 4 and 5)
@@
Table 4: Differential Demorphing: We compute the True Match Rate (TMR in %) at various False Match Rate (FMR in %) thresholds between the generated output (OUT) and the ground truth image (GT) across the five FR models used.
@@
Figure 4: Reference-Free Demorphing Visualization: We illustrate the results of our reference-free demorphing approach. The model takes the morph image (MORPH) as input and produces two outputs (OUT1 and OUT2). The ground truth images used to generate the morph are denoted as BF1 and BF2.
@@ 5.2.1 Reference Free Demorphing
We visualize the outputs of our reference-free demorphing method across six morphing techniques in Figure 4. These techniques include both landmark-based and deep-learning-based morphs. The first three columns visualize the morph and the two constituent face images used to create it, while the subsequent two columns show the outputs of our method, presented in no specific order. We observe that the faces produced by our method are visually distinct from each other and also appear notably different from the original morph. Moreover, they are visually closer to their ground-truth counterparts (refer to Appendix for more images). To quantify this, we calculate the normalized biometric distance between the generated faces and between the generated faces and the ground truth. We use five well-known face biometric models and plot the distance scores across three different pairings of images: OUT1-OUT2, BF1-OUT1, and BF2-OUT2, which represent the dissimilarity between the generated outputs and between each generated output and its corresponding ground truth. We plot the distance distribution for all datasets across all FR models in Figure 6. We observe an overall OUT1-OUT2 biometric distance of 0.4380, indicating a quantifiable difference in the output faces. Note that we use cosine distance rather than similarity to quantify results. This choice helps us intuitively explain morph-replication because a greater distance indicates that the generated faces are more distinguishable from each other. For BF1-OUT1, we observe a distance of 0.3992, and for BF2-OUT2, a distance of 0.4025. Note that the distance scores between outputs and their respective ground-truth face images fall well below the threshold used in several biometric systems [37, 36]. Our method successfully learns to distinguish features from the base image, such as hair (Rows 1 and 3), skin color, and subtle details like eyebrows and facial hair. Finally, we compute the True Match Rate (TMR) of the output subjects at various False Match Rate (FMR) thresholds and present the results in Table 3. Our method achieves an average TMR of 93.86% on the AMSL dataset. The corresponding TMR values on the OpenCV, StyleGAN, WebMorph, FaceMorph, and MorDiff datasets are 93.99%, 58.92%, 89.87%, 94.39%, and 93.75%, respectively. We also evaluate the quality of images generated by our method using various image-quality assessment (IQA) metrics, as shown in Table 6. Our method achieves an average FID [17] of 0.4284 across all datasets, with corresponding averages of 0.61 for SSIM [41], 0.24 for LPIPS [43], and 18.02 for PSNR, demonstrating the visual realism of the outputs from our method. We also observe that ArcFace performs consistently across all datasets in the presence of visual artefacts generated during demorphing compared to other FRS which is consistent with [5].
@@
Effect of $\mathcal{L}_{cr}$ and $\mathcal{E}(\cdot)$: We conduct an ablation study on our method to assess i) the impact of cross-road loss and ii) the influence of dual conditioning on generation. Excluding pixel-wise guidance (the second part of the loss in Equation 9) results in an average performance decline of 20.87% across the six datasets tested on the ArcFace FRS. This decrease is mainly due to the model’s reduced ability to accurately align the two unordered generated outputs with the ordered ground-truth images. Moreover, when the generator is trained without $\mathcal{E}$, the performance drops by 8.57% due to lack of context during the generation (see Section 3). These results are presented in Table 9.